OpenStudy (anonymous):

if a square had an area of 400m squared then what would its perimeter be

OpenStudy (anonymous):



OpenStudy (anonymous):

area will be equal to side^2 implies s^2=400 "s=20" now perimeter is given as 4*s=(s+s+s+s) implies perimeter=80

OpenStudy (anonymous):

@hamishawesome What do you think? Do you understand how to do a problem like this now?